OpenStudy (anonymous):

assume a general exponential growth/decay model. A certain population doubles every 15.75 years. How long will it take for it to triple?

OpenStudy (anonymous):

a. 24.96 years b. 23.625 years c. 21.854 years d. 25.15 years

ganeshie8 (ganeshie8):

\(2 = e^{15.75r}\) => \(r = \frac{\ln2}{15.75}\) ---------(1) since we need to find time it takes to triple :- \(3 = e^{rt}\) => \(t = \frac{\ln 3}{r}\) substitute r value above and see what u get

OpenStudy (perl):